# Optimizing Estimators of Squared Calibration Errors in Classification

### v1

@@ Abstract
In this work, we propose a mean-squared error-based risk that enables the comparison and optimization of estimators of squared calibration errors in practical settings. Improving the calibration of classifiers is crucial for enhancing the trustworthiness and interpretability of machine learning models, especially in sensitive decision-making scenarios. Although various calibration (error) estimators exist in the current literature, there is a lack of guidance on selecting the appropriate estimator and tuning its hyperparameters. By leveraging the bilinear structure of squared calibration errors, we reformulate calibration estimation as a regression problem with independent and identically distributed (i.i.d.) input pairs. This reformulation allows us to quantify the performance of different estimators even for the most challenging calibration criterion, known as canonical calibration. Our approach advocates for a training-validation-testing pipeline when estimating a calibration error on an evaluation dataset. We demonstrate the effectiveness of our pipeline by optimizing existing calibration estimators and comparing them with novel kernel ridge regression-based estimators on standard image classification tasks.
@@ 2 Background
In the following, we offer an extensive introduction into the background of this work. First, we give briefly measure theoretic preliminaries, followed by the different notions of calibration used for classification by the literature. We then introduce and discuss commonly used estimators for these notions.
@@ 2.1 Measure Theoretic Preliminaries
Throughout this work we use measure theoretic definitions to formalise our contribution, its assumptions, and its limitations. Specifically, we assume a measure space $\left(\Omega,\mathscr{F},\mu\right)$ is given. We associate with a random variable $X\colon\Omega\to\mathscr{X}$ (a measureable function) a probability space $\left(\mathscr{X},\mathscr{F}_{X},\mathbb{P}_{X}\right)$ with $\sigma$-field $\mathscr{F}_{X}=\left\{X\left(A\right)\mid A\in\mathscr{F}\right\}$ and pushforward measure $\mathbb{P}_{X}=\mu\circ X^{-1}$. Further, we may also assume a second random variable $Y\colon\Omega\to\mathscr{Y}$ such that we have a joint probability space $\left(\mathscr{X}\times\mathscr{Y},\mathscr{F}_{XY},\mathbb{P}_{XY}\right)$, where $\mathscr{F}_{XY}$ and $\mathbb{P}_{XY}$ are defined analogous as $\mathscr{F}_{X}$ and $\mathbb{P}_{X}$. Further, if a random variable $Y$ is discrete, we use $\mathbb{P}_{Y}$ to represent both its probability measure and the associated probability vector in the simplex $\Delta^{d}\coloneqq\left\{\left(p_{1},\dots,p_{k}\right)^{\intercal}\in\left[0,1\right]^{d}\mid\sum_{i=1}^{d}p_{i}=1\right\}$, where $d$ is the number of unique outcomes. In the same sense we may use $\mathbb{P}_{Y\mid X}\coloneqq\frac{\mathrm{d}\mathbb{P}_{XY}}{\mathrm{d}\mathbb{P}_{X}}$ as a measurable function $\mathscr{X}\to\Delta^{d}$. Finally, we say a statement holds $\mu$-almost surely ($\mu$-a.s.) if it is true except on a set of $\mu$-measure zero (Capiński & Kopp, 2004).
@@ 2.2 Mean-Squared Error Risk Minimization and Calibration
$\operatorname{CCE}_{2}\left(f\right)\coloneqq\sqrt{\mathbb{E}\left[\left\lVert f\left(X\right)-\mathbb{P}_{Y\mid f\left(X\right)}\right\rVert^{2}\right]}\;,$ (7)
@@ 2.3 Alternative Notions of Calibration
Besides canonical calibration, multiple notions of calibration have been introduced in the literature (Zadrozny & Elkan, 2002; Vaicenavicius et al., 2019; Kull et al., 2019; Gupta & Ramdas, 2022). Respective calibration errors assess the degree a classifier violates a given notion. In recent literature, the most common notion is top-label confidence calibration (Naeini et al., 2015; Guo et al., 2017; Joo et al., 2020; Kristiadi et al., 2020; Rahimi et al., 2020; Tomani et al., 2021; Minderer et al., 2021; Tian et al., 2021; Islam et al., 2021; Menon et al., 2021; Morales-Álvarez et al., 2021; Gupta et al., 2021; Wang et al., 2021; Fan et al., 2022; Dehghani et al., 2023; Chang et al., 2024). Here, we compare if the predicted top-label confidence $\max_{i\in\mathscr{Y}}f_{i}\left(X\right)$ matches the conditional accuracy $\mathbb{P}\left(Y=\arg\max_{i\in\mathscr{Y}}f_{i}\left(X\right)\mid\max_{i\in\mathscr{Y}}f_{i}\left(X\right)\right)$ given the prediction. Formally, we say the classifier $f$ is top-label confidence calibrated if and only if
@@
$\operatorname{TCE}_{2}\left(f\right)\coloneqq\sqrt{\mathbb{E}\left[\left(\max_{i}f_{i}\left(X\right)-\mathbb{P}\left(Y=\operatorname*{arg\,max}_{i}f_{i}\left(X\right)\mid\max_{i}f_{i}\left(X\right)\right)\right)^{2}\right]}\;.$ (9)
@@ 2.4 Calibration Estimators
$\begin{split}&\operatorname{TCE}_{2}^{\operatorname{bin}}\coloneqq\sqrt{\sum_{m=1}^{M}\frac{\lvert B_{m}\rvert}{n}\left(\operatorname{acc}\left(B_{m}\right)-\operatorname{conf}\left(B_{m}\right)\right)^{2}}\end{split}$ (10)
@@
$\begin{split}&\operatorname{CCE}_{2}^{\operatorname{kde}}\coloneqq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left\lVert f\left(X_{i}\right)-\frac{\sum_{j}k_{\operatorname{dir}}\left(f\left(X_{j}\right);f\left(X_{i}\right)\right)e_{Y_{j}}}{\sum_{j}k_{\operatorname{dir}}\left(f\left(X_{j}\right);f\left(X_{i}\right)\right)}\right\rVert^{2}},\end{split}$ (11)
where $e_{i}$ refers to the unit vector with a $1$ at index $i$ and $k_{\operatorname{dir}}$ is chosen to be the Dirichlet kernel, which is specifically suited for the simplex space (Ouimet & Tolosana-Delgado, 2022). The authors also propose analogous kernel density based estimators for top-label and class-wise calibration errors, which we will denote as $\operatorname{TCE}_{2}^{\operatorname{kde}}$ and $\operatorname{CWCE}_{2}^{\operatorname{kde}}$. Even though the kernel density approach is advantageous compared to the binning approach for higher dimensions, it still shares some of its limitations. Popordanoska et al. (2022b) show that the estimator converges in the infinite data limit, but it is still not clear what is the optimal choice of kernel and kernel hyperparameters in the finite data regime.
@@ 3.1 Theoretical Definition and Properties
$\mathbb{E}_{X}\left[h^{*}\left(f\left(X\right),f\left(X\right)\right)\right]=\operatorname{\operatorname{CCE}_{2}}.$ (13)
@@
This property would be sufficient in practice, if we were using $h$ with arguments sampled from $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$ to estimate the calibration error. However, as demonstrated by Equation (13), we predict $\operatorname{CCE}_{2}$ via the same sample in both arguments. Mirroring the arguments results in a possible $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$-null set since only elements in the diagonal $\mathscr{D}\left(\Delta^{d}\right)\coloneqq\left\{\left(p,p\right)\mid p\in\Delta^{d}\right\}\subset\Delta^{d}\times\Delta^{d}$ are considered. Consequently, the diagonal of an optimum $h^{\prime*}$ identified via $\mathscr{R}_{\operatorname{CE}}$ may "slip through" the $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$-a.s. guarantee in Theorem 1, and in turn may result in $\mathbb{E}\left[h^{\prime*}\left(f\left(X\right),f\left(X\right)\right)\right]\neq\operatorname{\operatorname{CCE}_{2}}$. To avoid such theoretical exceptions, we require additional theoretical assumptions, which we state in the following.
@@
$\mathscr{R}_{\operatorname{CE}}\left(h\right)=\mathscr{R}_{\operatorname{CE}}\left(h^{*}\right)\implies\mathbb{E}\left[h\left(f\left(X\right),f\left(X\right)\right)\right]=\operatorname{\operatorname{CCE}_{2}}.$ (19)
@@ 3.2.2 Calibration-Evaluation Pipeline
In general, we cannot expect to find $h^{*}$. However, the empirical risk allows us to find a parametrized $h_{\theta,\eta}$ close to $h^{*}$, where $\theta$ denote its parameters and $\eta$ its hyperparameters. Consequently, similar to how traditional machine learning works, we need a training-validation-test split to achieve unbiased estimation of the calibration error once we optimized $h_{\theta,\eta}$. Specifically, we use a training set to find $\theta_{\operatorname{tr}}$ and a validation set to find $\eta_{\operatorname{val}}$. The final calibration estimate is then computed by
$\widehat{\operatorname{CE}}_{2}\left(f\right)\coloneqq\frac{1}{n_{\operatorname{te}}}\sum_{X\in\mathscr{D}_{\operatorname{te}}}h_{\theta_{\operatorname{tr}},\eta_{\operatorname{val}}}\left(f\left(X\right),f\left(X\right)\right),$ (21)
where $\mathscr{D}_{\operatorname{te}}$ is the test set of size $n_{\operatorname{te}}$, and $\widehat{\operatorname{CE}}_{2}$ is representative for different notions of calibration. We might also use multiple training, validation and test sets via (nested) cross validation. In that case, we average the results of the calibration estimation functions fitted in each fold.
@@
We encourage to never compare the test set risk of different calibration estimation functions, since this would put a bias on the final calibration estimation. Neglecting this is equivalent to selecting an optimal classifier based on the test accuracy in a classification task.
4 Calibration Estimation Functions
@@
4.1 Binning and Kernel Density
@@
for some $h\colon\Delta^{d}\times\Delta^{d}\to\mathbb{R}$ and an i.i.d. dataset $\left(X_{1},Y_{1}\right),\dots,\left(X_{n},Y_{n}\right)\sim\mathbb{P}_{XY}$. For the binning based estimator, define
@@
$\left(\operatorname{TCE}_{2}^{\mathrm{bin}}\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{bin}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right).$ (24)
@@
$\left(\operatorname{CCE}_{2}^{\mathrm{kde}}\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{kde}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right).$ (26)
We will also use an analogous estimator $\operatorname{TCE}_{2}^{\mathrm{kde}}$ for estimating $\operatorname{TCE}_{2}$ in the experiment sections. Extending the binning based estimator to $\operatorname{CCE}_{2}$ is in general not possible. Note that the runtime complexity of $\operatorname{CCE}_{2}^{\mathrm{kde}}$ is in $O\left(n^{2}d\right)$, since we compare every evaluation instance with every training instance for every class. In the following we introduce novel estimators, which are runtime invariant to the number of classes. They are based on kernel ridge regression and directly minimize the empirical risk under kernel ridge regression assumptions.
4.2 Kernel Ridge Regression

### v2

@@ Abstract
In this work, we propose a mean-squared error-based risk that enables the comparison and optimization of estimators of squared calibration errors in practical settings. Improving the calibration of classifiers is crucial for enhancing the trustworthiness and interpretability of machine learning models, especially in sensitive decision-making scenarios. Although various calibration (error) estimators exist in the current literature, there is a lack of guidance on selecting the appropriate estimator and tuning its hyperparameters. By leveraging the bilinear structure of squared calibration errors, we reformulate calibration estimation as a regression problem with independent and identically distributed (i.i.d.) input pairs. This reformulation allows us to quantify the performance of different estimators even for the most challenging calibration criterion, known as canonical calibration. Our approach advocates for a training-validation-testing pipeline when estimating a calibration error on an evaluation dataset. We demonstrate the effectiveness of our pipeline by optimizing existing calibration estimators and comparing them with novel kernel ridge regression-based estimators on real-world image classification tasks.
@@ 2 Background
In the following, we offer an extensive introduction to the background of this work. First, we offer a more extensive motivation for calibration. Second, we give briefly measure theoretic preliminaries, followed by the different notions of calibration used for classification by the literature. We then introduce and discuss commonly used estimators for these notions.
2.1 An Exemplary Motivation for Calibration
[…]
@@ 2.2 Measure Theoretic Preliminaries
Throughout this work we use measure theoretic definitions to formalize our contribution, its assumptions, and its limitations. Specifically, we assume a measure space $\left(\Omega,\mathscr{F},\mu\right)$ is given. We associate with a random variable $X\colon\Omega\to\mathscr{X}$ (a measureable function) a probability space $\left(\mathscr{X},\mathscr{F}_{X},\mathbb{P}_{X}\right)$ with $\sigma$-field $\mathscr{F}_{X}=\left\{X\left(A\right)\mid A\in\mathscr{F}\right\}$ and pushforward measure $\mathbb{P}_{X}=\mu\circ X^{-1}$. Further, we may also assume a second random variable $Y\colon\Omega\to\mathscr{Y}$ such that we have a joint probability space $\left(\mathscr{X}\times\mathscr{Y},\mathscr{F}_{XY},\mathbb{P}_{XY}\right)$, where $\mathscr{F}_{XY}$ and $\mathbb{P}_{XY}$ are defined analogous as $\mathscr{F}_{X}$ and $\mathbb{P}_{X}$. Further, if a random variable $Y$ is discrete, we use $\mathbb{P}_{Y}$ to represent both its probability measure and the associated probability vector in the simplex $\Delta^{d}\coloneqq\left\{\left(p_{1},\dots,p_{k}\right)^{\intercal}\in\left[0,1\right]^{d}\mid\sum_{i=1}^{d}p_{i}=1\right\}$, where $d$ is the number of unique outcomes. In the same sense we may use $\mathbb{P}_{Y\mid X}\coloneqq\frac{\mathrm{d}\mathbb{P}_{XY}}{\mathrm{d}\mathbb{P}_{X}}$ as a measurable function $\mathscr{X}\to\Delta^{d}$. Finally, we say a statement holds $\mu$-almost surely ($\mu$-a.s.) if it is true except on a set of $\mu$-measure zero (Capiński & Kopp, 2004).
@@ 2.3 Mean-Squared Error Risk Minimization and Calibration
$\operatorname{CCE}_{2}\left(f\right)\coloneqq\sqrt{\mathbb{E}_{X}\left[\left\lVert f\left(X\right)-\mathbb{P}_{Y\mid f\left(X\right)}\right\rVert^{2}\right]}\;,$ (7)
@@ 2.4 Alternative Notions of Calibration
Besides canonical calibration, multiple notions of calibration have been introduced in the literature (Zadrozny & Elkan, 2002; Vaicenavicius et al., 2019; Kull et al., 2019; Gupta & Ramdas, 2022). Respective calibration errors assess the degree to which a classifier violates a given notion. In recent literature, the most common notion is top-label confidence calibration (Naeini et al., 2015; Guo et al., 2017; Joo et al., 2020; Kristiadi et al., 2020; Rahimi et al., 2020; Tomani et al., 2021; Minderer et al., 2021; Tian et al., 2021; Islam et al., 2021; Menon et al., 2021; Morales-Álvarez et al., 2021; Gupta et al., 2021; Wang et al., 2021; Fan et al., 2022; Dehghani et al., 2023; Chang et al., 2024). Here, we compare if the predicted top-label confidence $\max_{i\in\mathscr{Y}}f_{i}\left(X\right)$ matches the conditional accuracy $\mathbb{P}\left(Y=\arg\max_{i\in\mathscr{Y}}f_{i}\left(X\right)\mid\max_{i\in\mathscr{Y}}f_{i}\left(X\right)\right)$ given the prediction. Formally, we say the classifier $f$ is top-label confidence calibrated if and only if
@@
$\operatorname{TCE}_{2}\left(f\right)\coloneqq\sqrt{\mathbb{E}_{X}\left[\left(\max_{i}f_{i}\left(X\right)-\mathbb{P}\left(Y=\operatorname*{arg\,max}_{i}f_{i}\left(X\right)\mid\max_{i}f_{i}\left(X\right)\right)\right)^{2}\right]}\;.$ (9)
@@ 2.5 Calibration Estimators
$\begin{split}&\operatorname{TCE}_{2}^{\operatorname{bin}}\left(f\right)\coloneqq\sqrt{\sum_{m=1}^{M}\frac{\lvert B_{m}\rvert}{n}\left(\operatorname{acc}\left(B_{m}\right)-\operatorname{conf}\left(B_{m}\right)\right)^{2}}\end{split}$ (10)
@@
$\begin{split}&\operatorname{CCE}_{2}^{\operatorname{kde}}\left(f\right)\coloneqq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left\lVert f\left(X_{i}\right)-\frac{\sum_{j}k_{\operatorname{dir}}\left(f\left(X_{j}\right);f\left(X_{i}\right)\right)e_{Y_{j}}}{\sum_{j}k_{\operatorname{dir}}\left(f\left(X_{j}\right);f\left(X_{i}\right)\right)}\right\rVert^{2}},\end{split}$ (11)
where $e_{i}$ refers to the unit vector with a $1$ at index $i$ and $k_{\operatorname{dir}}$ is chosen to be the Dirichlet kernel, which is specifically suited for the simplex space (Ouimet & Tolosana-Delgado, 2022). The authors also propose analogous kernel density based estimators for top-label and class-wise calibration errors, which we will denote as $\operatorname{TCE}_{2}^{\operatorname{kde}}$ and $\operatorname{CWCE}_{2}^{\operatorname{kde}}$. Even though the kernel density approach is advantageous compared to the binning approach for higher dimensions, it still shares some of its limitations. Popordanoska et al. (2022b) show that the estimator converges in the infinite data limit, but it is still not clear what is the optimal choice of kernel and kernel hyperparameters in the finite data regime. Further, the runtime complexity $\operatorname{CCE}_{2}^{\mathrm{kde}}$ is in $O\left(n^{2}\right)$.
@@ 3.1 Theoretical Definition and Properties
$\sqrt{\mathbb{E}_{X}\left[h^{*}\left(f\left(X\right),f\left(X\right)\right)\right]}=\operatorname{\operatorname{CCE}}_{2}\left(f\right).$ (13)
@@
This property would be sufficient in practice, if we were using $h$ with arguments sampled from $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$ to estimate the calibration error. However, as demonstrated by Equation (13), we predict $\operatorname{CCE}_{2}$ via the same sample in both arguments. Mirroring the arguments results in a possible $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$-null set since only elements in the diagonal $\mathscr{D}\left(\Delta^{d}\right)\coloneqq\left\{\left(p,p\right)\mid p\in\Delta^{d}\right\}\subset\Delta^{d}\times\Delta^{d}$ are considered. Consequently, the diagonal of an optimum $h^{\prime*}$ identified via $\mathscr{R}_{\operatorname{CE}}$ may "slip through" the $\mathbb{P}_{f\left(X\right)}\otimes\mathbb{P}_{f\left(X\right)}$-a.s. guarantee in Theorem 1, and in turn may result in $\sqrt{\mathbb{E}\left[h^{\prime*}\left(f\left(X\right),f\left(X\right)\right)\right]}\neq\operatorname{CCE}_{2}\left(f\right)$. To avoid such theoretical exceptions, we require additional theoretical assumptions, which we state in the following.
@@
$\mathscr{R}_{\operatorname{CE}}\left(h\right)=\mathscr{R}_{\operatorname{CE}}\left(h^{*}\right)\implies\sqrt{\mathbb{E}_{X}\left[h\left(f\left(X\right),f\left(X\right)\right)\right]}=\operatorname{\operatorname{CCE}}_{2}\left(f\right).$ (19)
@@ 3.2.2 Calibration-Evaluation Pipeline
In general, we cannot expect to find $h^{*}$. However, the empirical risk allows us to find a $\hat{h}_{\eta}\in H_{\eta}$ close to $h^{*}$, where $H_{\eta}$ is a model class with hyperparameter $\eta\in\Theta$ and search space $\Theta$. Examples for $\eta$ are the number of bins in Equation (10) or the bandwidth in Equation (11). Consequently, similar to traditional machine learning, we need a training-validation-test split to achieve a debiased estimation of the calibration error once we optimized $\eta$. Specifically, we use a training set to fit $\hat{h}_{\eta}$ and a validation set to find an optimal $\eta_{\operatorname{val}}$. The final calibration estimate is then computed by
$\widehat{\operatorname{CE}}_{2}\left(f\right)\coloneqq\sqrt{\frac{1}{n_{\operatorname{te}}}\sum_{X\in D_{\operatorname{te}}}\hat{h}_{\eta_{\operatorname{val}}}\left(f\left(X\right),f\left(X\right)\right)},$ (21)
where $D_{\operatorname{te}}$ is the test set of size $n_{\operatorname{te}}$, and $\widehat{\operatorname{CE}}_{2}$ is representative for different notions of calibration. We may also use multiple training and validation sets via cross-validation, similar to conventional hyperparameter optimization (Bischl et al., 2023). Our proposed pipeline for estimating the calibration error of a classifier via cross-validation is presented in Algorithm 1. The algorithm mimics hyperparameter optimization of a conventional machine learning model (Bischl et al., 2023). Consequently, using our pipeline also implies a computational complexity of $O\left(k\lvert\Theta\rvert\right)$, where $k$ is the number of folds in cross-validation. In comparison, using a default hyperparameter, as in (Guo et al., 2017) or (Minderer et al., 2021), has $O(1)$ complexity since no optimization happens.
@@
We recommend not comparing the holdout test set risk of different calibration estimation functions since this would put a bias on the final calibration estimation. Neglecting this is similar to selecting an optimal classifier based on the test accuracy in a classification task.
Algorithm 1 Evaluating the calibration of a given classifier and dataset by optimizing the calibration estimator. The evaluation dataset is split into a holdout set for estimating the calibration error, and another set, which is used for optimizing the calibration estimator via cross-validation.
  Input: Evaluation dataset $D=\left\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\right\}$, classifier $f$, model class $H_{\eta}$ of calibration estimation functions with hyperparameter search space $\Theta\ni\eta$, $k$ number of folds.
  $D_{\mathrm{pr}}\leftarrow\left\{(f\left(X_{1}\right),Y_{1}),\dots,(f\left(X_{n}\right),Y_{n})\right\}\quad$ > compute classifier predictions
  $D_{\mathrm{opt}},D_{\mathrm{te}}\leftarrow\operatorname{split}\left(D_{\mathrm{pr}}\right)\quad$ > create optimization and holdout test set
  $\{(D_{1}^{\mathrm{train}},D_{1}^{\mathrm{eval}}),\dots,(D_{k}^{\mathrm{train}},D_{k}^{\mathrm{eval}})\}\leftarrow\operatorname{CVfolds}\left(D_{\mathrm{opt}}\right)\quad$ > create cross-validation folds
  for $i=1,\dots,k$ do
     for $\eta\in\Theta$ do
        $\hat{h}_{\eta}^{i}\leftarrow\operatorname{fit}\left(H_{\eta},D_{i}^{\mathrm{train}}\right)\quad$ > fit for a given hyperparameter and training set
        $\mathrm{risk}_{\eta}^{i}\leftarrow\hat{\mathscr{R}}_{\mathrm{CE}}\left(\hat{h}_{\eta}^{i}\right)\quad$ > compute risk for $\hat{h}_{\eta}$ with data $D_{i}^{\mathrm{eval}}$ according to Eq. (20)
     end for
  end for
  $\eta_{\operatorname{val}}\leftarrow\arg\min_{\eta\in H}\frac{1}{k}\sum_{i=1}^{k}\mathrm{risk}_{\eta}^{i}\quad$ > get the hyperparameter with the smallest average risk
  Define $\hat{h}_{\eta_{\operatorname{val}}}\left(x,y\right)\coloneqq\frac{1}{k}\sum_{i=1}^{k}\hat{h}_{\eta_{\operatorname{val}}}^{i}\left(x,y\right)$
  Compute $\widehat{\operatorname{CE}}_{2}\left(f\right)$ with $\hat{h}_{\eta_{\operatorname{val}}}$ and $D_{\mathrm{te}}$ according to Eq. (21)
  return  $\widehat{\operatorname{CE}}_{2}\left(f\right)$
In this section, we have established our optimization and evaluation pipeline based on calibration estimation functions. Next, we show how already existing calibration error estimators used in the literature can be framed as calibration estimation functions.
4 Calibration Error Estimators as Calibration Estimation Functions
@@
4.1 Reframing Binning and Kernel Density based Calibration Estimators
@@
for some $h\colon\Delta^{d}\times\Delta^{d}\to\mathbb{R}$ and an i.i.d. dataset $\left(X_{1},Y_{1}\right),\dots,\left(X_{n},Y_{n}\right)\sim\mathbb{P}_{XY}$. For the binning-based estimator, define
@@
$\left(\operatorname{TCE}_{2}^{\mathrm{bin}}\left(f\right)\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{bin}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right).$ (24)
@@
$\left(\operatorname{CCE}_{2}^{\mathrm{kde}}\left(f\right)\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{kde}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right).$ (26)
We will also use an analogous estimator $\operatorname{TCE}_{2}^{\mathrm{kde}}$ for estimating $\operatorname{TCE}_{2}$ in the experiment sections. Extending the binning-based estimator to $\operatorname{CCE}_{2}$ is in general not possible. Further, our calibration evaluation pipeline allows us to compare the risk of different calibration estimators. This motivates the introduction of novel calibration estimators, especially for canonical calibration errors, to have a larger search space to optimize over. In the following, we introduce a novel class of calibration estimators based on kernel ridge regression, which minimizes the empirical risk in Equation (20) plus a regularization term under typical kernel ridge regression assumptions.
4.2 Novel Calibration Estimators based on Kernel Ridge Regression